Let P(A) = 0.6, P(B) = 0.3, and P(ABC) = 0.1. Calculate P(A|B).
Answer #1

P(A Bc ) = P(A) + P(B) - P(A Bc )

0.1 = 0.6 + 0.3 - P(A Bc )

P(A Bc ) = 0.8

Now

P(A Bc ) = P(A) - P(A B)

0.8 = 0.6 - P(A B)  

P(A B) = 0.2

Therefore

P(A \ B) = P(A B) / P(B)

= 0.2 / 0.3

= 2/3

= 0.6667